OpenStudy (anonymous):

Part 1: Explain, in complete sentences, which method you would use to solve the following system of equations. (1 point) Part 2: Explain why you chose that method (1 point) Part 3: Provide the solution to the system. (2 points) x – 3y + 2z = –12 x + 2y + 3z = 6 2x – 3y – z = –2

OpenStudy (anonymous):

Part 1 - Substitution. Part 2 - since You can use the value of 1 variable in the terms of the other 2 variables and then form any 2 equations in 2 variables and solve accordingly. Part 3 - x-3y+2z = -12 Hence x = 3y-2z-12 Use x = 3y-2z-12 in x+2y+3z = 6 , and 2x-3y-z = -2 You will get two equations having only Y, Z as the variables. Solve for Y,Z hence solve for x and post the answer. :)